Instant hot packs work by crystallizing sodium acetate (NaCH3COO). The molar enthalpy of crystallization for sodium acetate is -56.7 kJ/mol. How many grams of sodium acetate are needed to warm 125.0 mL of water from 21.0 oC to 35.4oC?


1
Expert's answer
2021-08-30T02:25:14-0400


The molar enthalpy of crystallization for sodium acetate is -56.7 kJ/mol


And latent heat of vaporisation of water is 40.662 KJ/mol


Amount of heat required to raise the temperature of 125 ml of H2O from 21°C to 35.4°C


Density of water is = 1 g/ml

That means we have 125 g of H2O

OR

125/18=6.94 mole of water

Total amount heat to raise temperature is =nC∆t

=6.94×40.662×14.4

=4066.2 KJ/mol

No of mole of sodium acetate required is

=4066.2 / 56.7

=71.71 mole of sodium acetate

Mass of sodium acetate required

=No. Of mole sodium acetate×molecular Mass of sodium acetate

=71.71×82

=5880.6 g of sodium acetate